# Spin-orbit interaction in tubular prismatic nanowires

### v1

@@ Abstract
We theoretically study the spin-orbit interaction in the outer regions of core-shell nanowires that can act as tubular, prismatic conductors. The polygonal cross section of these wires induces non-uniform electron localization along the wire perimeter. In particular, low-energy electrons accumulate in the corner regions, and in the case of narrow shells, conductive channels form along the sharp edges. In contrast, higher-energy electrons are shifted toward the facets. These two groups of states may be separated by large energy gaps, which can exceed the room-temperature energy in the case of triangular geometries. We compare the impact of spin-orbit interaction on the corner and side states of hexagonal and triangular shells grown on hexagonal cores as well as on triangular shells grown on triangular cores. We find that the spin-orbit splitting, and thus the degeneracy of energy states at finite wave vectors, strongly depends on the tube’s geometry. We demonstrate that the weak spin-orbit coupling observed in clean wires can be significantly enhanced if the intermixing of core and shell materials takes place. Moreover, we show that the energy spectrum in the presence of spin-orbit interaction allows for estimating the interaction between states and shows that triangular shells can act as three independent wires in the low-energy regime, while they behave as interacting systems at higher energy ranges.
@@ I Introduction
Core-(multi-)shell nanowires are radial heterojunctions formed by overgrowing a quantum wire with layers of different materials. Although cylindrical structures have been obtained [14], in most cases the crystallographic properties of the components lead to polygonal cross sections. The most common are hexagonal nanowires [15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25], but triangular [26, 27, 28, 29, 30], rectangular [31, 32], and even dodecagonal [33] shapes have also been demonstrated. Furthermore, nanowires combining two different polygons within a single cross section are technologically feasible. Specifically, triangular shells can be grown on hexagonal cores such that either the vertices of the internal and external boundaries match [34], or their sides are parallel [35].
@@
The combination of two semiconductors with significantly different band gaps can result in a type I band alignment at the junction between the materials, leading to the transfer of electrons to the material with the narrower gap. By properly adjusting the material and geometric parameters, it becomes possible to create a radial heterojunction in which electrons are confined in the core region [36, 37, 20] or in the shell region [15, 18, 25, 19]. In terms of electron localization, the former case resembles single-material wires, where the electron ground state is localized at the center of the sample [36, 38]. As a result, the shape of the cross section does not significantly affect the properties of the electron in its lowest-energy state. For higher energy states, the probability distributions exhibit several maxima, which are determined by the number of vertices. However, these maxima often overlap, causing electrons excited to higher energies to remain near the center of the cross-section.
As a result, the properties of electrons confined in the cores of core-shell wires or in single-material wires, and in particular the SOI, can be effectively described using a cylindrical wire model [39]. However, this description does not apply when the electrons occupy the shell region. The electron distribution in the shell differs significantly between the circular and polygonal cross sections. In cylindrical wires, the electrons are uniformly distributed along the perimeter of the cross section. In contrast, for prismatic tubes, the electrons tend to localize in different regions of the cross section, depending on their energy.
@@
In the present paper, we consider core-shell nanowires with hexagonal and triangular cross sections and calculate the effects of SOI arising from the extended band offset potential and an external electric field on electronic energy spectra using the $\bm{k}\!\cdot\!\bm{p}$ method. We show that the SOI-induced degeneracy differs between the hexagonal shell and the other geometries studied. Moreover, the effect of SOI is significantly enhanced in structures where intermixing of the core and shell materials has occurred. In the presence of SOI, the corner states of triangular shells resemble the spectrum of three independent wires, while the lowest side-localized states exhibit clear signatures of interaction.
@@
The paper is organized as follows. Section II introduces the model, while Section III presents the results for different geometries. Conclusions are summarized in Section IV.
@@ II Model
We study the electrons confined in the shells of an infinitely long InP-InAs core-shell nanowires with three different cross-sectional geometries: a triangular shell grown on a triangular or hexagonal core (Tri-Hex), and a hexagonal shell covering a hexagonal core, as shown in Fig. 1. Our computational approach can be applied to core–shell nanowires with shells that have the zinc-blende crystallographic structure and can be easily modified to describe structures with wurtzite shells. We chose InP and InAs materials because both Tri-Hex [35] and hexagonal [40] InP–InAs core–shell nanowires have been demonstrated. Moreover, triangular InP nanowires [41], which could be used as cores, have also been experimentally realized.
Figure 1: Cross sections of hexagonal (a), triangular (b) and Tri-Hex (c) wires. $R_{\mathrm{c}}$ and $R_{\mathrm{s}}$ represent the external radii of the core and shell (wire), respectively. The shell thickness is denoted by $d$ and the intermixing length is indicated by $r$. $R_{\mathrm{c}}=40$ nm, $d=8$ nm, and $r=2$ nm.
We consider wires grown along the crystallographic direction [111], for which the Dresselhaus SOI can be neglected [42]. The 8-band Kane Hamiltonian is given by
@@
where $H_{cc}$ and $H_{vv}$ are diagonal matrices $2\times 2$ and $6\times 6$, respectively, describing the two lowest conduction bands and the six highest valence bands. Assuming the wires are grown along the $z$-axis, the conduction-band Hamiltonian is:
@@
includes the contributions due to conduction-band offset which penetrates into the shell [$V_{\mathrm{BO}}(x,y)$] and the potential due to the external electric field [$V_{\mathrm{E}}(x,y)$].
We assume that initially the intermixing of the core and shell materials, InP and InAs, takes place. However, the proportion of core material gradually decreases while that of the shell material increases during the shell growth process [43]. In our model, the volume in which the contributions from both materials are present is defined by its cross section, i.e. a ring surrounding the core, darker blue regions in Fig. 1. The thickness of this ring is denoted by r [Fig. 1(b)] and referred to as the intermixing length. In general, r can range from zero to the arbitrary value, but in this study, we assume a minimal intermixing length of $r=2$ nm. This controlled material mixing results in a tilted potential within the shell, which corresponds to an extended band offset potential ($V_{\mathrm{BO}}$) that (in our model) decreases linearly along the radial lines of the grid from a value of $V_{\mathrm{CBO}}=656.6$ meV [44] at the core-shell boundary, which is the conduction band offset at the InP-InAs heterojunction.
If the distance $r$ is less than the thickness of the shell $d$ (Fig. 1), the potential drops to zero, i.e. reaches the bottom of the shell material conduction band at the distance $r$ from the core-shell interface. If $r\geqslant d$ the entire shell contains InP molecules and the value of potential $V_{\mathrm{BO}}$ at the wire external boundary increases with $r$, but stays within the interval $(0,V_{\mathrm{CBO}})$. This potential gradient creates a built-in, or intrinsic, radial electric field in the wire, thereby inducing SOI. Furthermore, we consider the possibility of an external electric field ($V_{\mathrm{E}}$) applied perpendicular to the wire, which can alter the electron distribution within the shell [45] and further influence the SOI.
@@
$H_{vv}=\mathrm{diag}\left[\begin{array}[]{c}V(x,y)-E_{g},\\ V(x,y)-E_{g},\\ V(x,y)-E_{g},\\ V(x,y)-E_{g},\\ V(x,y)-E_{g}-\Delta_{so},\\ V(x,y)-E_{g}-\Delta_{so}\end{array}\right]\,,$ (13)
where $E_{g}$ is the semiconductor band gap and $\Delta_{so}$ is the spin-split gap.
The off-diagonal block is:
@@
is proportional to the non-vanishing momentum matrix elements between the conduction ($|S\sigma\rangle$) and the valence $(|X(Y,Z)\sigma\rangle$) band-edge Bloch states at the $\Gamma$-point of the Brillouin zone. Here, $\sigma$ represents spin and $p_{\pm}=p_{x}\pm ip_{y}$.
The folding-down transformation [46, 47] allows us to obtain the reduced Hamiltonian that describes only the conduction electrons, but with the influence of the highest valence bands on their energies. For the shell, this Hamiltonian takes the form:
@@
is the effective mass, and the Rashba spin-orbit coefficients are:
@@
is the amplitude of the Rashba coefficients, and $\sigma_{x}$ and $\sigma_{y}$ are the Pauli matrices [48, 49, 50]. The expressions for the effective mass [Eq. (20)] and the Rashba coefficients’ amplitude [Eq. (22)] were derived based on the assumption that the semiconductor band gap of the shell ($E_{g}=417$ meV) is the largest energy within the studied subsystem. Note that for the narrow shells, the values of $m^{*}$ and $\alpha_{0}$ (Eqs. (20) and (22), respectively) might get slightly modified - but this effect is beyond our present approximation since our goal is to establish the qualitative picture of the SOI in core-shell nanowires, in particular to specify the degeneracy of the states.
To obtain the energy dispersion, we use a discretization method developed in Ref. 51. We construct a polar grid, on which we superimpose polygonal constraints corresponding to the cross-sectional geometry of the shell, and diagonalize the Hamiltonian [Eq. (II)] for a large set of $k_{z}$ values, one at a time, on the reduced grid located within the shell. We apply Dirichlet and the infinite potential well boundary conditions at the shell-vacuum and core-shell polygonal boundaries [45, 52, 25]. To compute the transverse states, we numerically diagonalize the Hamiltonian in Eq. (II) for $k_{z}=0$ using the effective mass of the conduction band [Eq. (20)] [45, 52, 25]. This method allows us to model wires with arbitrary cross-sectional shapes, including non-symmetric cases [45, 53].
Our approach includes the geometrical confinement of the electrons inside the narrow tubular shells, but we neglect the effects of the Coulomb repulsion between electrons, typically represented by the Hartree potential, to be obtained with the Poisson equation. Because we consider a low concentration of electrons confined in a narrow shell of 6-18 nm thickness, the redistribution of electrons due to electrostatic forces is minimal, but computationally demanding. Nevertheless, we tested a couple of situations and indeed the corrections to the energy spectra due to the Hartree potential were not significant.
@@ III Results
The numerical simulations presented below were performed for cross sections consisting of over 6,000 grid points for hexagonal (Hex) shells, over 7,000 for triangular (Tri) wires, and between 6,200 and 7,200 for Tri-Hex structures. To preserve cross-sectional symmetry, the number of angular grid points was chosen to be a multiple of the number of internal vertices. The material parameters were either taken directly from or calculated based on Ref. 44. For the InAs shell, these parameters are: $E_{g}=417$ meV, $\Delta_{so}=390$ meV, and $P_{0}=9$ Js. For InP, the band gap is $E_{g}=1423.6$ meV. The InP-InAs valence band offset is $350$ meV, and the resulting conduction band offset is $V_{\mathrm{CBO}}=656.6$ meV.
Figure 2: Probability distributions corresponding to the lowest energy level (lower row) and the first energy level above the gap $\Delta$ (upper row) for hexagonal (a), triangular (b), and Tri-Hex (c) cross sections ($d=8$ nm, $R_{\mathrm{c}}=40$ nm).
Figure 3: Low-energy transverse states for the cross sections shown in Fig. 1. The gap $\Delta$ separating the lowest (corner) states is: $\Delta_{\mathrm{Hex}}=6$ meV, $\Delta_{\mathrm{Tri}}=33$ meV, and $\Delta_{\mathrm{Tri-Hex}}=31$ meV.
Figure 4: Low-energy states for the cross sections shown in Fig. 1 ($d=8$ nm, $R_{\mathrm{c}}=40$ nm) without SOI.
@@ III.1 Transverse states
In contrast to circular rings, electrons in polygonal rings are non-uniformly distributed along the perimeter. Specifically, low-energy electrons tend to localize in the corner regions, where they are attracted by effective quantum wells, like in the case of bent wires [54]. The curved region of a bent wire behaves like a quantum well, whose depth and length are determined by the bend’s angle and radius. A physically analogous effect arises in the corner regions of polygonal quantum rings. As a result, the probability distributions exhibit maxima in these areas (Fig. 3, lower row). For hexagonal cross sections, the maxima spread widely and overlap along the sides [Fig. 3(a), lower panel], allowing the carriers to move freely around the cross section. However, in triangular and Tri-Hex structures, the electrons with the lowest energies are depleted from the sides [Figs. 3(b) and 3(c), lower row]. Here, the probability distributions form sharp peaks localized at the corners, leading to the formation of well-separated conductive channels. This means that a single shell may contain multiple 1D wires, each confined to one of the sharp edges [55]. When the shell is confined internally and externally by the same polygon, higher transverse states are localized along the sides [Figs. 3(a) and 3(b), upper row] [56, 45]. In the Tri-Hex structure, the higher states remain localized at the corners, but the probability maxima form two peaks which are slightly shifted towards the sides [Fig. 3(c), upper row].
@@
In Fig. 3, we compare the transverse energy states for the three cross sections shown in Fig. 1. The energy states are either 2-fold or 4-fold degenerate, due to spin only or spin and geometric symmetry, respectively, and are grouped according to their localization characteristics. The number of states in each group is twice the number of external vertices (Fig. 3). The dispersion of the lowest, corner-localized states decreases with the number of external vertices and side thickness [45, 52]. For sufficiently narrow triangular and Tri-Hex shells, the corner states become quasi-degenerate [Figs. 3(b) and 3(c)]. If the shells were thicker, the lowest states would split into a 2-fold degenerate ground state and a 4-fold degenerate excited state [52].
@@
The corner states are energetically separated from the higher states by an energy gap $\Delta$, which increases with decreasing side thickness and the number of corners [45, 52]. The quantum wells associated with the corners are deep when the ratio between the effective linear size of the corner area and the thickness of the polygon sides is large.
@@
For the hexagonal shell, the gap is larger than the energy differences within the corner-localized states but is comparable to the gaps between higher levels. In triangular and Tri-Hex shapes, the gap separating the six lowest states is significantly larger than other energy splittings and can even exceed the room-temperature energy. For the geometry and material parameters considered here, $\Delta_{\mathrm{Tri}}$ is slightly larger than $\Delta_{\mathrm{Tri-Hex}}$, but the case, where $\Delta_{\mathrm{Tri-Hex}}$ is considerably larger than $\Delta_{\mathrm{Tri}}$ is also possible [57]. Moreover, proper adjustment of the shell’s cross-section shape and material properties may result in an energy separation of the corner states exceeding 100 meV [45, 57]. Although the energy gap $\Delta$ decreases with increasing shell thickness [52], for triangular and Tri-Hex structures it remains considerably large even for very wide shells.
@@
Thus, prismatic shells provide a subspace of corner-localized states that are expected to be robust against perturbations, such as weak disorder or cross-sectional symmetry breaking.

### v2

@@ Abstract
We theoretically study the spin-orbit interaction in the outer regions of core-shell nanowires that can act as tubular, prismatic conductors. The polygonal cross section of these wires induces non-uniform electron localization along the wire perimeter. In particular, low-energy electrons accumulate in the corner regions, and in the case of narrow shells, conductive channels form along the sharp edges. In contrast, higher-energy electrons are shifted toward the facets. These two groups of states may be separated by large energy gaps, which can exceed the room-temperature energy in the case of triangular geometries. We compare the impact of spin-orbit interaction on the corner and side states of hexagonal and triangular shells grown on hexagonal cores as well as on triangular shells grown on triangular cores. We find that the spin-orbit splitting, and thus the degeneracy of energy states at finite wave vectors, strongly depend on the tube’s geometry. We demonstrate that the weak spin-orbit coupling observed in clean wires can be significantly enhanced if the intermixing of core and shell materials takes place. Moreover, we show that the energy spectrum in the presence of spin-orbit interaction allows for estimating the interaction between states and shows that triangular shells can act as three independent wires in the low-energy regime, while they behave as interacting systems at higher-energy ranges.
@@ I Introduction
Core-shell nanowires are radial heterojunctions formed by overgrowing a quantum wire with layers of different materials. Although cylindrical structures have been obtained [14], in most cases the crystallographic properties of the components lead to polygonal cross sections. The most common are hexagonal nanowires [15, 16, 17, 18, 19, 20, 21, 22, 23, 24, 25], but triangular [26, 27, 28, 29, 30], rectangular [31, 32], and even dodecagonal [33] shapes have also been demonstrated. Furthermore, nanowires combining two different polygons within a single cross section are technologically feasible. Specifically, triangular shells can be grown on hexagonal cores such that either the vertices of the internal and external boundaries match [34], or their sides are parallel [35].
@@
The combination of two semiconductors with significantly different band gaps can result in a type I band alignment at the junction between the materials, leading to the transfer of electrons to the material with the narrower gap. By properly adjusting the material and geometric parameters, it becomes possible to create a radial heterojunction in which electrons are confined in the core region [36, 37, 20] or in the shell region [15, 18, 25, 19]. In terms of electron localization, the former case resembles single-material wires, where the electron ground state is localized at the center of the sample [36, 38]. As a result, the shape of the cross section does not significantly affect the properties of the electron in its lowest-energy state. For higher-energy states, the probability distributions exhibit several maxima, which are determined by the number of vertices. However, these maxima often overlap, causing electrons excited to higher energies to remain near the center of the cross section. As a result, the properties of electrons confined in the cores of core-shell wires or in single-material wires, and in particular the SOI, can be effectively described using a cylindrical wire model [39]. However, this description does not apply when the electrons occupy the shell region. The electron distribution in the shell differs significantly between the circular and polygonal cross sections. In cylindrical wires, the electrons are uniformly distributed along the circumference of the cross section. In contrast, for prismatic tubes, the electrons tend to localize in different regions of the cross section, depending on their energy.
@@
In the present paper, we consider core-shell nanowires with hexagonal and triangular cross sections and calculate the effects of SOI arising from the extended band-offset potential and an external electric field on electronic energy spectra using the $\bm{k}\!\cdot\!\bm{p}$ method. We show that the SOI-induced degeneracy differs between the hexagonal shell and the other geometries studied. Moreover, the effect of SOI is significantly enhanced in structures where intermixing of the core and shell materials occurs. In the presence of SOI, the corner states of triangular shells resemble the spectrum of three independent wires, while the lowest side-localized states exhibit clear signatures of interaction.
@@
The paper is organized as follows. Section II introduces the model, while Sec. III presents the results for different geometries. Conclusions are summarized in Sec. IV.
@@ II Model
We study the electrons confined in the shells of infinitely long InP-InAs core-shell nanowires with three different cross-sectional geometries: a triangular shell grown on a triangular or hexagonal core (Tri-Hex), and a hexagonal shell covering a hexagonal core, as shown in Fig. 1. Our computational approach can be applied to core-shell nanowires with shells that have the zinc-blende crystallographic structure and can be easily modified to describe structures with wurtzite shells. We chose InP and InAs materials because both Tri-Hex [35] and hexagonal [40] InP-InAs core-shell nanowires have been demonstrated. Moreover, triangular InP nanowires [41], which could be used as cores, have also been experimentally realized.
Figure 1: Cross sections of (a) hexagonal, (b) triangular, and (c) Tri-Hex wires. $R_{\mathrm{c}}$ and $R_{\mathrm{s}}$ represent the external radii of the core and shell (wire), respectively. The shell thickness is denoted by $d$ and the intermixing length is indicated by $r$. $R_{\mathrm{c}}=40$ nm, $d=8$ nm, and $r=2$ nm.
We consider wires grown along the crystallographic direction [111], for which the Dresselhaus SOI can be neglected [42]. The eight-band Kane Hamiltonian is given by
@@
where $H_{cc}$ and $H_{vv}$ are diagonal matrices $2\times 2$ and $6\times 6$, respectively, describing the two lowest conduction bands and the six highest valence bands. Assuming the wires are grown along the $z$ axis, the conduction-band Hamiltonian is
@@
includes the contributions due to conduction-band offset which penetrates into the shell $V_{\mathrm{BO}}(x,y)$] and the potential due to the external electric field $V_{\mathrm{E}}(x,y)$].
We assume that initially the intermixing of the core and shell materials, InP and InAs, takes place. However, the proportion of core material gradually decreases while that of the shell material increases during the shell growth process [43]. In our model, the volume in which the contributions from both materials are present is defined by its cross section, i.e., a ring surrounding the core (darker blue regions in Fig. 1). The thickness of this ring is denoted by $r$ [Fig. 1(b)] and referred to as the intermixing length. In general, $r$ can range from zero to the arbitrary value, but in this study, we assume a minimal intermixing length of $r=2$ nm. This controlled material mixing results in a tilted potential within the shell, which corresponds to an extended band-offset potential ($V_{\mathrm{BO}}$) that (in our model) decreases linearly along the radial lines of the grid from a value of $V_{\mathrm{CBO}}=656.6$ meV [44] at the core-shell boundary, which is the conduction-band offset at the InP-InAs heterojunction.
If the distance $r$ is less than the thickness of the shell ($d$ in Fig. 1), the potential $V_{\mathrm{BO}}$ drops to zero, i.e., reaches the bottom of the shell material conduction band at the external boundary of the polygon defined by $r$, i.e., within the shell, and the outer region of the shell consists solely of InAs. If $r\geqslant d$ the entire shell contains P atoms and the value of potential $V_{\mathrm{BO}}$ at the wire external boundary increases with $r$, but stays within the interval $(0,V_{\mathrm{CBO}})$. This potential gradient creates a built-in, or intrinsic, radial electric field in the wire, thereby inducing SOI. Furthermore, we consider the possibility of an external electric field ($V_{\mathrm{E}}$) applied perpendicular to the wire, which can alter the electron distribution within the shell [45] and further influence the SOI.
@@
$H_{vv}=\mathrm{diag}\left[\begin{array}[]{c}V(x,y)-E_{g},\\ V(x,y)-E_{g},\\ V(x,y)-E_{g},\\ V(x,y)-E_{g},\\ V(x,y)-E_{g}-\Delta_{\mathrm{so}},\\ V(x,y)-E_{g}-\Delta_{\mathrm{so}}\end{array}\right]\,$, (13)
where $E_{g}$ is the semiconductor band gap and $\Delta_{\mathrm{so}}$ is the spin-split gap.
The off-diagonal block is
@@
is proportional to the nonvanishing momentum matrix elements between the conduction ($|S\sigma\rangle$) and the valence $(|X(Y,Z)\sigma\rangle$) band-edge Bloch states at the $\Gamma$ point of the Brillouin zone. Here, $\sigma$ represents spin and $p_{\pm}=p_{x}\pm ip_{y}$.
The folding-down transformation [46, 47] allows us to obtain the reduced Hamiltonian that describes only the conduction electrons, but with the influence of the highest valence bands on their energies. For the shell, this Hamiltonian takes the form
@@
is the effective mass, and the Rashba spin-orbit coefficients are
@@
is the amplitude of the Rashba coefficients, and $\sigma_{x}$ and $\sigma_{y}$ are the Pauli matrices [48, 49, 50]. The expressions for the effective mass [Eq. (20)] and the Rashba coefficients’ amplitude [Eq. (22)] were derived based on the assumption that the semiconductor band gap of the shell ($E_{g}=417$ meV) is the largest energy within the studied subsystem. Note that for the narrow shells, the values of $m^{*}$ and $\alpha_{0}$ [Eqs. (20) and (22), respectively] might get slightly modified—but this effect is beyond our present approximation since our goal is to establish the qualitative picture of the SOI in core-shell nanowires, in particular to specify the degeneracy of the states.
To obtain the energy dispersion, we use a discretization method developed in Ref. 51. We construct a polar grid, on which we superimpose polygonal constraints corresponding to the cross-sectional geometry of the shell, and diagonalize the Hamiltonian [Eq. (II)] for a large set of $k_{z}$ values, one at a time, on the reduced grid located within the shell. We apply the Dirichlet and the infinite potential well boundary conditions at the shell-vacuum and core-shell polygonal boundaries [45, 52, 25]. To compute the transverse states, we numerically diagonalize the Hamiltonian in Eq. (II) for $k_{z}=0$ using the effective mass of the conduction band [Eq. (20)] [45, 52, 25]. This method allows us to model wires with arbitrary cross-sectional shapes, including nonsymmetric cases [45, 53].
Our approach includes the geometrical confinement of the electrons inside the narrow tubular shells, but we neglect the effects of the Coulomb repulsion between electrons, typically represented by the Hartree potential, to be obtained with the Poisson equation. Because we consider a low concentration of electrons confined in a narrow shell of 8–24 nm thickness, the redistribution of electrons due to electrostatic forces is minimal, but computationally demanding. Nevertheless, we tested a couple of situations and indeed the corrections to the energy spectra due to the Hartree potential were not significant.
@@ III Results
The numerical simulations presented below were performed for cross sections consisting of over 6000 grid points for hexagonal shells, over 7000 for triangular wires, and between 6200 and 7200 for Tri-Hex structures. To preserve cross-sectional symmetry, the number of angular grid points was chosen to be a multiple of the number of internal vertices. The material parameters were either taken directly from or calculated based on Ref. 44. For the InAs shell, these parameters are: $E_{g}=417$ meV, $\Delta_{\mathrm{so}}=390$ meV, and $P_{0}=9\times 10^{-10}$ eVs. For InP, the band gap is $E_{g}=1423.6$ meV. The InP-InAs valence-band offset is $350$ meV, and the resulting conduction-band offset is $V_{\mathrm{CBO}}=656.6$ meV.
Figure 2: Probability distributions corresponding to the lowest energy level (lower row) and the first energy level above the gap $\Delta$ (upper row) for (a) hexagonal, (b) triangular, (c) and Tri-Hex cross sections ($d=8$ nm, $R_{\mathrm{c}}=40$ nm).
Figure 3: Low-energy transverse states for the cross sections shown in Fig. 1. The gap $\Delta$ separating the lowest (corner) states is: $\Delta_{\mathrm{Hex}}=6$ meV, $\Delta_{\mathrm{Tri}}=33$ meV, and $\Delta_{\mathrm{Tri\textrm{-}Hex}}=31$ meV.
Figure 4: Low-energy states for the cross sections shown in Fig. 1 without SOI. The energies are shifted with respect to the ground state energy at $k_{z}=0$. (a) States $l=1$, $l=4$, $l=5$, and $l=8$ are twofold degenerate, while states $l=2$, $l=3$, $l=6$, and $l=7$ are fourfold degenerate. (b) and (c) States $l=1$ and $l=4$ are twofold degenerate, while states $l=2$ and $l=3$ are fourfold degenerate.
@@ III.1 Transverse states
In contrast to circular rings, electrons in polygonal rings are nonuniformly distributed along the perimeter. Specifically, low-energy electrons tend to localize in the corner regions, where they are attracted by effective quantum wells, as in the case of bent wires [54]. The curved region of a bent wire behaves as a quantum well, whose depth and length are determined by the bend’s angle and radius. A physically analogous effect arises in the corner regions of polygonal quantum rings. As a result, the probability distributions exhibit maxima in these areas (Fig. 3, lower row). For hexagonal cross sections, the maxima spread widely and overlap along the sides [Fig. 3(a), lower panel], allowing the carriers to move around the cross section. However, in triangular and Tri-Hex structures, the electrons with the lowest energies are depleted from the sides [Figs. 3(b) and 3(c), lower row]. Here, the probability distributions form sharp peaks localized at the corners, leading to the formation of well-separated conductive channels. This means that a single shell may contain multiple one-dimensional wires, each confined to one of the sharp edges [55]. When the shell is confined internally and externally by the same polygon, higher transverse states are localized along the sides [Figs. 3(a) and 3(b), upper row] [56, 45]. In the Tri-Hex structure, the higher states remain localized at the corners, but the probability maxima form two peaks in each corner [Fig. 3(c), upper row].
@@
In Fig. 3, we compare the transverse energy states for the three cross sections shown in Fig. 1. The energy states are either twofold or fourfold degenerate, due to spin only or spin and geometric symmetry, respectively, and are grouped according to their localization characteristics. The number of states in each group is twice the number of external vertices (Fig. 3). The dispersion of the lowest, corner-localized states decreases with the number of external vertices and side thickness [45, 52]. For sufficiently narrow triangular and Tri-Hex shells, the corner states become quasidegenerate [Figs. 3(b) and 3(c)]. If the shells were thicker, the lowest states would visibly split into a twofold degenerate ground state and a fourfold degenerate excited state [52].
@@
The effective quantum wells associated with the corners deepen when the ratio between the effective linear size of the corner area and the thickness of the polygon sides increases, thus they become deeper when the number of corners decreases. As a result, the energy gap $\Delta$, which separates the corner states, increases with decreasing the number of corners and also side thickness [45, 52].
@@
For the 8-nm-wide hexagonal shell, the gap is larger than the energy differences within the corner-localized states but is comparable to the gaps between higher levels. In triangular and Tri-Hex shapes, the gap separating the six lowest states is significantly larger than other energy splittings and can even exceed the room-temperature energy. For the geometry and material parameters considered here, $\Delta_{\mathrm{Tri}}$ is slightly larger than $\Delta_{\mathrm{Tri\textrm{-}Hex}}$, but the case, where $\Delta_{\mathrm{Tri\textrm{-}Hex}}$ is considerably larger than $\Delta_{\mathrm{Tri}}$ is also possible [57]. Moreover, a proper adjustment of the shell’s cross-section shape and material properties may result in an energy separation of the corner states exceeding 100 meV [45, 57]. Although the energy gap $\Delta$ decreases with increasing shell thickness [52], for triangular and Tri-Hex structures it remains considerably large even for very wide shells.
@@
Thus, prismatic shells provide a well-separated subspace of corner-localized states that are expected to be robust against perturbations, such as weak disorder or cross-sectional symmetry breaking.